OpenStudy (anonymous):

Help please for equation that represents the line that passes through the point (-2, 6) and has a slope of 4?

jimthompson5910 (jim_thompson5910):

Start with y = mx+b where m is the slope and b is the y-intercept

jimthompson5910 (jim_thompson5910):

You're given a slope of 4. So m = 4 One point that lies on the line is (-2,6). So x = -2 and y = 6 So y = mx+b becomes 6 = (4)(-2)+b Now solve for b

OpenStudy (anonymous):

How do I do that....

jimthompson5910 (jim_thompson5910):

6 = (4)(-2)+b will become 6 = -8+b what's next?

OpenStudy (anonymous):

@jim_thompson5910 do I subtract 8?

jimthompson5910 (jim_thompson5910):

6 = -8+b is the same as 6 = b-8 so what's next?

OpenStudy (anonymous):

divide?

jimthompson5910 (jim_thompson5910):

add 8 to both sides

OpenStudy (anonymous):

so 14 = b :D

jimthompson5910 (jim_thompson5910):

yes, which means that y = mx+b becomes y = 4x+14

jimthompson5910 (jim_thompson5910):

since m = 4 and b = 14

OpenStudy (anonymous):

Oh okay thanks

jimthompson5910 (jim_thompson5910):

yw